Jessica has two children: Sophie and Lucas. Jessica is six times as old as Sophie, who is two years older than Lucas. In six years, Jessica will be four times as old as Lucas. How old is Jessica?
Found 2 solutions by ankor@dixie-net.com, iammewearethree:
Answer by ankor@dixie-net.com(22740)

Jessica has two children: Sophie and Lucas.
Let j = Jessica's age
Let s = Sophie's age
Let L = Lucas's ge
:
Write an equation for each statement
:
Jessica is six times as old as Sophie,
j = 6s
who is two years older than Lucas.
s = L + 2
or
L = s - 2
In six years, Jessica will be four times as old as Lucas.
j + 6 = 4(L + 6)
Replace j with 6s
Replace L with (s-2)
6s + 6 = 4((s-2)+6)
6s + 6 = 4(s + 4)
6s + 6 = 4s + 16
6s - 4s = 16 - 6
2s = 10
s = 5 yrs is Sophie's age
:
How old is Jessica? She's 6 times as old as Sophie
6 * 5 = 30 yrs

Answer by iammewearethree(1)

Jessica is 30 years old. Jessica=6s Sophie=2+L Lucas=S-2 6s+6=4((s-2)+6). Solve for s and you will get 30.